a number increased by twice the square root is 48. what is this number
Answer by checkley71(8403)

X+2SQRTX=48
2SQRTX=48-X
NOW SQUARE BOTH SIDES OF THE EQUATION.
4X=2304-96X+X^2
X^2-96X-4X+2304=0
X^2-100X+2304=0
(X-36)(X-64)=0
X-36=0
X=36 ANSWER.
PROOF
36+2SQRT36=48
36+2*6=48
36+12=48
48=48